# Conflict-Based Search as a Protocol: A Multi-Agent Motion Planning Protocol for Heterogeneous Agents, Solvers, and Independent Tasks

### v1

@@ IV-A Single-Agent Solvers
A* Edge validity Successor Generator [30] Ackermann, 4-connected, Dubins Start-goal, Coverage
@@
Diffusion Penalty Warm start [26] None Start-goal, Motion patterns
@@ IV-A1 Heuristic Search
We implement a single-agent A* search as a representative heuristic search algorithm. At its core, heuristic search methods build a graph of different motion primitives and searches over them with the help of a heuristic. We plan in continuous space for robots with dynamic constraints (i.e., Ackermann and Dubins) by using motion primitives. We use a modified version of soft-duplicate detection [36] to speed up search.
@@
Experience: We implement experience by using the technique in [30]. When we find a solution, we store the nodes on the solution path. Then when replanning, if we encounter a node in the solution path, we add as successors the rest of the nodes in the path that are still valid under the new constraints.
@@ IV-A2 Sampling Based Search
We implement Rapidly-exploring Random Tree (RRT) [37] as a representative of sampling based search algorithms. We chose to use a regular spatial RRT for its simplicity and popularity (with each vertex having an associated time of arrival). We note this means that the RRT solution does not contain wait actions and that the solver is incomplete. However, it can still be used in a plan() API (it may just return failure when solutions exist).
@@
Experience: Given a new query, we iterate through all nodes/edges in the tree created in the previous query and remove those invalid under the new constraints. We then start the RRT process from this tree. We note more sophisticated versions of replanning using experience exist (e.g., [38]).
@@ IV-A3 Optimization
We formulate the motion planning problem as a trajectory optimization problem, whose objective is to drive the system from the initial state to the goal in minimum time. To this end, we transcribe the problem into a nonlinear program, with a cost function that minimizes a weighted objective and constraints that enforce both dynamic feasibility and spatial validity. Dynamic constraints are implemented using direct collocation [39], while collision avoidance is addressed through a duality-based exact representation [40]. We initialize the nonconvex nonlinear program with an A* solution and solve it using IPOPT [41].
@@ IV-A4 Diffusion
We implement a diffusion solver by modifying Multi-robot Multi-model planning Diffusion (MMD) [26] for our map and tasks.
@@ IV-A5 Reinforcement Learning
We implement an Reinforcement Learning (RL) solver using PPO [22]. To our knowledge, we are unaware of existing papers that have used an RL single-agent solver with CBS. The model takes in a local observation and outputs a categorical probability distribution of the next action. Concretely, the model inputs are:
@@
    An estimate of distance to goal computed via a backward Dijkstra assuming the agent has gridworld dynamics (so not aware of the agent’s footprint and dynamics),
@@
The model then outputs a 1-hot probability distribution corresponding to a discrete set of motion primitives for a robot with Dubins dynamics. We utilize motion primitives to handle kinodynamic constraints and as it also enables us to use action masking to speed up training. To generate a full path, we rollout the policy, greedily following its actions and backtracking if necessary when hitting a deadend (i.e., DFSing based on the agent’s policy).
@@ IV-B CBS Protocol Settings
We primarily care about the success rate (as opposed to solution cost) in our experiments and therefore use a “Greedy” CBS variant [42] for the CBS Protocol that prioritizes CT nodes by number of pairs of agents with conflicts and tiebreaks by solution cost. We use discrete collision checking every 0.1 simulation seconds to detect conflicts. Each robot has a fixed footprint, so the APIs output space-time volumes by outputting their space-time path and footprint, which is then used by CBS for collision detection. Constraints for a collision at $(x_{0},y_{0},t_{0})$ block a spatial square of size 0.1 containing $x_{0},y_{0}$ for 2.5 seconds. These numbers were semi-arbitrarily picked for performance and to emphasize continuous space / continuous time constraints and planning.
@@ IV-C Experimental Results
We test our method on the Cantwell (7.2x15.2, Fig 2), Ribera (8.7x6.3), and Mosquito (23.4x11.1) maps from the IN2D dataset [43] that realistically model a house or office space, as well as an empty map (10x15) and large (32x32) map with 10% obstacles (size in meters). Each robot has a footprint between 0.1 to 0.8 meters (some circles, rectangles, and triangles). All single-agent APIs/solvers, collision-checking, and the CBS Protocol were written in Python and experiments were run on a machine with an i7-11800H@2.30GHz x 16 CPU and 3050 Ti Laptop GPU.
@@
Recall from Section III-A that the initial “root node” created in CBS requires planning each agent individually without constraints. Thus, the number of “root conflicts” serves as an approximation for problem difficulty as having more conflicts implies a more congested problem instance that requires more work for the CBS Protocol to resolve. Additionally, a root node with 0 conflicts indicates that individually planning and executing would result in a successful instance. Of the 675 instances, 175 instances fell into this category.
Fig 3 shows the success rate (bottom) and number of CT Nodes generated (top) based on the number of root conflicts. The CBS Protocol solved 54% of instances with non-zero conflicts. As the number of conflicts increases, the number of CT Nodes generated for successful instances increases (i.e., the min and median of the blue dots increases). We also observe that the overall success rate decreases as root conflicts increase.
@@
Heterogeneous Tasks: We do a proof-of-concept and replace the start-goal task in the plan API with different tasks. For A* and RRT we implement coverage planners that need to visit a set of points in any arbitrary order. For optimization (direct collocation), the agent attempts to maintain a radius around a given “surveillance” point. For diffusion, we train a policy to mimic motion patterns on data where the robot visits an intermediate point before proceeding to the goal. Finally, for the RL solver we implement a Roomba-like policy that tries to sweep the entire environment. Due to space-constraints, Figure 2 shows a single example solution with various APIs (more examples provided in the video). We highlight that we just need to input the different plan APIs into the CBS Protocol without any additional changes.
Figure 3: We show results on MAMP start-goal problems with different solvers and categorize them by the number of conflicts in the root node (a proxy for problem difficulty). The top plot shows the number of CT nodes that the CBS Protocol generates for successful (blue) and unsuccessful (red) instances while the bottom plot shows the overall success rate.
@@ V Conclusion and Future Directions
The near future of robotics will require multiple robots from different companies completing individual tasks to efficiently find collision-free paths in shared environments. It is unclear how to effectively do this due to each robot having its own interfaces and potential proprietary motion planning software.
To the end, we show how we can use Conflict-Based Search [2] as a protocol. The CBS Protocol solely requires agents to specify a single plan() API that takes in certain space-time constraints and returns a path with corresponding cost and space-time volume. We show how the CBS Protocol enables finding collision-free paths for a variety of solvers(A*, RRT, Direct Collocation, Diffusion, Reinforcement Learning), and is not limited to start-goal tasks but can work more broadly with arbitrary independent motion planning tasks.
Future Work: We believe that the CBS Protocol and AH-MAMP perspective open promising avenues for future research. From an industry perspective, we would like to see the CBS Protocol adopted and deployed on real heterogeneous systems with efficient solvers.
From the single-agent plan() API perspective, there is significant potential in developing better solvers that handle CBS space-time constraints and replanning. In particular, these space-time constraints are a special subcategory of “dynamic” obstacles whose structure could be exploited for efficient planning. More effective versions of reusing experience in respect to these constraints is also promising work.
From the AH-MAMP and multi-agent protocol perspective, the use of heterogeneous plan() APIs opens up interesting questions on how to best leverage their strengths and weaknesses. For example, in our experiments, different APIs had varying success rates and query times. Enabling the CBS Protocol to intelligently reason about this would likely boost performance. Additionally, as a proof of concept, the current version of the CBS Protocol generalizes the core ideas of CBS from classical MAPF but does not include additional improvements in the MAPF literature. We envision bringing more enhancements (e.g., better space-time constraints, an ECBS version [42], interleaving planning and execution [44]) as important steps for increasing the practicality of the CBS Protocol. Achieving this would likely require expanding the single-agent plan() API inputs.
@@
To our knowledge, the problem of multi-agent motion planning (MAMP) with algorithmically heterogeneous single-agent solvers has not been addressed by the community. Current heterogeneous MAMP methods are not algorithmically heterogeneous as they require the different robots to all run the same planner [16, 11] or focus on heterogeneous behaviors (and not collision-free movement) [27, 28]. We introduce and formalize this new setting as Algorithmically Heterogeneous MAMP (AH-MAMP). Solutions for AH-MAMP cannot rely on accessing the internals of each single-agent solver. Instead, they must be formulated as protocols that interact with single-agent planners only through well-defined APIs.
@@
Our second contribution is to cast Conflict-Based Search (CBS) [2]—a foundational method in multi-agent path finding [3]—as such a protocol. We observe that CBS fundamentally relies on a single-agent plan() interface that can return a path subject to space-time constraints. Specifically, the API requires:
@@
As a proof-of-concept, the CBS Protocol took the core CBS idea without leveraging modern advancements in classical CBS. Thus, promising future work would be to incorporate improvements like stronger constraints and Enhanced CBS [42]. Interleaving partial planning and execution [44] could also replace the need for a single long planning time in the beginning with interspersed planning throughout. Achieving this would likely require expanding the single-agent plan() API inputs.

### v2

@@ IV-A Single-Agent Solvers
A* Edge validity Successor Generator [shaoul2024experience-cbs] Ackermann, 4-connected, Dubins Start-goal, Coverage
@@
Diffusion Penalty Warm start [shaoul2025mmd] None Start-goal, Motion patterns
@@ IV-A1 Heuristic Search
We implement a single-agent A* search as a representative heuristic search algorithm. We plan in continuous space for robots with dynamic constraints (i.e., Ackermann and Dubins) by using discretized motion primitives. We use a modified version of soft-duplicate detection [du2019soft_duplicate_detection] to speed up search.
@@
Experience: We implement experience by using the technique in [shaoul2024experience-cbs]. When we find a solution, we store the nodes on the solution path. Then when replanning, if we encounter a node in the solution path, we add as successors the rest of the nodes in the path that are valid under the new constraints.
@@ IV-A2 Sampling Based Search
We implement Rapidly-exploring Random Tree (RRT) [lavalle1998rrt] as a representative of sampling based search algorithms. We use a regular spatial RRT for its simplicity and popularity (with each vertex having an associated time of arrival). We note this means that the RRT solution does not contain wait actions and that the solver is incomplete. However, it can still be used in a plan() API (it may just return failure when solutions exist).
@@
Experience: Given a new query, we iterate through all nodes/edges in the tree created in the previous query and remove those invalid under the new constraints. We then start the RRT process from this tree. We note more sophisticated versions of replanning using experience exist (e.g., [ferguson2006replanning_with_rrts]).
@@ IV-A3 Optimization
We formulate the motion planning problem as a trajectory optimization problem, whose objective is to drive the system from the initial state to the goal in minimum time. To this end, we transcribe the problem into a nonlinear program, with a cost function that minimizes a weighted objective and constraints that enforce both dynamic feasibility and spatial validity. Dynamic constraints are implemented using direct collocation [kelly2017introduction], while collision avoidance is addressed through a duality-based exact representation [zhang2020optimization]. We initialize the nonconvex nonlinear program with an A* solution and solve it using IPOPT [wachter2006implementation].
@@ IV-A4 Diffusion
We implement a diffusion solver by modifying Multi-robot Multi-model planning Diffusion (MMD) [shaoul2025mmd] for our map and tasks.
@@ IV-A5 Reinforcement Learning
We implement an Reinforcement Learning (RL) solver using PPO [schulman2017ppo]. To our knowledge, we are unaware of existing papers that have used an RL single-agent solver with CBS. The model takes in a local observation and outputs a categorical probability distribution of the next action. Concretely, the model inputs are:
@@
    An estimate of distance to goal computed via a backward Dijkstra assuming gridworld dynamics (so unaware of the agent’s footprint or dynamics),
@@
The model outputs a 1-hot probability distribution corresponding to a discrete set of motion primitives for a robot with Dubins dynamics. We utilize motion primitives to handle kinodynamic constraints and as it also enables using action masking to speed up training. To generate a full path, we rollout the policy, greedily following its actions and backtracking if necessary when hitting a deadend (i.e., DFSing based on the agent’s policy).
@@ IV-B CBS Protocol Settings
We primarily care about the success rate (as opposed to solution cost) in our experiments and therefore use a “Greedy” CBS variant [barer2014suboptimalecbs] for the CBS Protocol that prioritizes CT nodes by number of pairs of agents with conflicts and tiebreaks by solution cost. We use discrete collision checking every 0.1 simulation seconds to detect conflicts. Each robot has a fixed footprint, so the APIs output space-time volumes by outputting their space-time path and footprint, which is then used by CBS for collision detection. Constraints for a collision at $(x_{0},y_{0},t_{0})$ block a spatial square of size 0.1 containing $x_{0},y_{0}$ for 2.5 seconds. These numbers were semi-arbitrarily picked for performance and to emphasize continuous space / continuous time constraints and planning.
@@ IV-C Experimental Results
We test our method on the Cantwell (7.2x15.2, Fig 2), Ribera (8.7x6.3), and Mosquito (23.4x11.1) maps from the IN2D dataset [dobrevskiskocaj2020in2d] that realistically model a house or office space, as well as an empty map (10x15) and large (32x32) map with 10% obstacles (size in meters). Each robot has a footprint between 0.1 to 0.8 meters (some circles, rectangles, and triangles). All single-agent APIs/solvers, collision-checking, and the CBS Protocol were written in Python and experiments were run on a machine with an i7-11800H@2.30GHz x 16 CPU and 3050 Ti Laptop GPU.
@@
Recall from Section III-A that the initial “root node” created in CBS requires planning each agent individually without constraints. Thus, the number of “root conflicts” (number of pairs of agents that conflicts) serves as an approximation for problem difficulty as having more conflicts implies a more congested problem instance that requires more work for the CBS Protocol to resolve. Additionally, a root node with 0 conflicts indicates that individually planning and executing would result in a successful instance. Of the 675 instances, 175 instances fell into this category.
Fig 3 shows the success rate (bottom) and number of CT Nodes generated (top) based on the number of root conflicts. The CBS Protocol solved 54% of instances with non-zero conflicts. As the number of conflicts increases, the number of CT Nodes generated for successful instances increases (e.g., the median of the blue dots increases). We also observe that the overall success rate decreases as root conflicts increase.
@@
Heterogeneous Tasks: We do a proof-of-concept and replace the start-goal task in the plan API with different tasks. For A* and RRT we implement coverage planners that need to visit a set of points in any arbitrary order. For optimization (direct collocation), the agent attempts to maintain a radius around a given “surveillance” point. For diffusion, we train a policy to mimic motion patterns on data where the robot visits an intermediate point before proceeding to the goal. Finally, for the RL solver we implement a Roomba-like policy that tries to sweep the entire environment. Due to space constraints, Figure 2 shows a single example solution with various APIs (more examples are provided on the website). We highlight that we only need to input the different plan APIs into the CBS Protocol without any additional changes.
Figure 3: We show results on MAMP start-goal problems with different solvers and categorize them by the number of conflicts in the root node (a proxy for problem difficulty). The top plot shows the number of CT nodes that the CBS Protocol generates for successful (blue) and unsuccessful (red) instances while the bottom shows the success rate.
@@ V Conclusion and Future Directions
The future of robotics will require multiple robots from different companies completing individual tasks to efficiently find collision-free paths in shared environments. It is unclear how to effectively do this due to each robot having its own interfaces and motion planning software. To that end, we show how we can use Conflict-Based Search [sharon2015cbs] as a protocol. The CBS Protocol solely requires agents to specify a single plan() API that takes in certain space-time constraints and returns a path with corresponding cost and space-time volume. We show how the CBS Protocol enables finding collision-free paths for a variety of solvers (A*, RRT, Direct Collocation, Diffusion, Reinforcement Learning), and is not limited to start-goal tasks but can work with arbitrary independent motion planning tasks.
Future Work: From an industry perspective, we would like to see the CBS Protocol adopted and deployed on real heterogeneous systems with efficient solvers.
From the single-agent plan() API perspective, there is significant potential in developing better solvers that handle CBS space-time constraints and replanning. In particular, these space-time constraints are a special subcategory of “dynamic” obstacles whose structure could be exploited for efficient planning. More effective versions of reusing experience in respect to these constraints is also promising.
From the AH-MAMP and multi-agent protocol perspective, the use of heterogeneous plan() APIs opens up interesting questions on how to best leverage their strengths and weaknesses. For example, in our experiments, different APIs had varying success rates and query times. Enabling the CBS Protocol to intelligently reason about this would likely boost performance. Additionally, as a proof of concept, the current version of the CBS Protocol retools CBS from classical MAPF but does not include additional improvements in the MAPF literature. Achieving this would likely require expanding the single-agent plan() API inputs.
@@
To our knowledge, the problem of multi-agent motion planning (MAMP) with algorithmically heterogeneous single-agent solvers has not been addressed by the community. Current heterogeneous MAMP methods are not algorithmically heterogeneous as they require the different robots to all run the same planner [debord2018multi_drone_ground_cbs, solis2021representation_optimal_cbs] or focus on heterogeneous behaviors (and not collision-free movement) [rizk2019heterogeneous_multi_robot_survey, bettini2023heterogeneous_behavior_marl]. We introduce and formalize this new setting as Algorithmically Heterogeneous MAMP (AH-MAMP). Solutions for AH-MAMP cannot rely on accessing the internals of each single-agent solver. Instead, they must be formulated as protocols that interact with single-agent planners only through well-defined APIs.
@@
Our second contribution is to cast Conflict-Based Search (CBS) [sharon2015cbs]—a foundational method in multi-agent path finding [stern2019mapfbenchmark]—as such a protocol. We observe that CBS fundamentally relies on a single-agent plan() interface that can return a path subject to space-time constraints. Specifically, the API requires:
@@
As a proof-of-concept, the CBS Protocol took the core CBS idea without leveraging modern advancements in classical CBS. Thus, promising future work would be to incorporate improvements like stronger constraints and Enhanced CBS [barer2014suboptimalecbs]. Interleaving partial planning and execution [rhcrLi2020] could also replace the need for a single long planning time in the beginning with interspersed planning throughout. Achieving this would likely require expanding the single-agent plan() API inputs.